OpenStudy (anonymous):

If the ball is released at a speed of 25 km/h from a height of 1.5 m, will the ball go over the shed? The polynomial function h(t) = -4.9t2 + 25t + 1.5 models this situation.

OpenStudy (anonymous):

25×1000 = 25000 m 1 hour = 60min × 60 sec/min= 3600 sec 1.5 m / (6.95 m/sec)= 0.216 sec h(t) = -4.9 (0.216)^2 +25 ( 0.216) +1.5 = 6.671 its seems to be true ... hope I'm right .. regards

OpenStudy (anonymous):

Thanks! It does make sense

OpenStudy (anonymous):

if forgot that 25 km/h = 6.95 m/sec ^_^!

OpenStudy (anonymous):

wait- can explain how did you 6.95? :3

OpenStudy (anonymous):

1 km = 1000 m this is fact 1 h = 60 min , its fact also 1 min = 60 sec ........ if you have 60 mints then you have to multiplying both sides by 60 any way 1km = 1000 m.... you have 25 km then you have to multiplying both sides by 25 25km/h =( 25000/3600)m/sec = 6.95 hope it clear ..

OpenStudy (anonymous):

ohhhhh okay! Thanks so much!!!